OpenStudy (anonymous):

I don't seem to understand how to approach this type of question: For the following function, find the direction in which the directional derivative is zero: w = xy + yz + xz, at (1,-1,2)

OpenStudy (anonymous):

I calculated the gradient at this point: \[grad(w) = <1,3,0>\] and now i think that I'm supposed to solve \[<x,y,z>*<1,3,0> = 0\] and: \[x^2+y^2+z^2 = 1\] but I'm lacking a 3rd equation for z... in the answer they just set z = c (for all c) and get some vector but it differs from mine and I don't get how they came to it... (my grad is correct though).

OpenStudy (anonymous):

Having multiple solutions is okay.

OpenStudy (anonymous):

Or are they expecting only one solution?

OpenStudy (anonymous):

that's their answer

OpenStudy (anonymous):

the top is -3i+j+ck

OpenStudy (anonymous):

Yes, there are infinite solutions, that why they have \(c\) as a free variable I suppose?

OpenStudy (anonymous):

yep, but I don't see how they came to this expression

OpenStudy (anonymous):

Well... \[ x+3y = 0 \]

OpenStudy (anteater):

What is their answer?

OpenStudy (anonymous):

Such is how they get the \(-3,1\)

OpenStudy (anonymous):

\(z\) becomes a free variable.

OpenStudy (anonymous):

However, I got this funky feeling that there may be more solutions than just that...

OpenStudy (anonymous):

Wait, never mind. They've got all the solutions.

OpenStudy (anonymous):

Well, thank you very much. I will play a bit more with the equations to see what can I get out of them. As far as I can see it's just algebric manipulation from now.